OpenStudy (anonymous):

if the sides of a square are lengthened by 8cm, the area becomes 169cm^2. Find the length of a side of the original square. The length of a side of the original square is ..... cm?

OpenStudy (anonymous):

169 is the square of 13

OpenStudy (anonymous):

13-8=5

OpenStudy (anonymous):

i didnt really understand the whole cm^2 thing to figure it out

OpenStudy (anonymous):

that just means area. area is given in terms of square units

OpenStudy (anonymous):

i say don't worry about it

OpenStudy (anonymous):

you wouldn't say the area of a table top is 10 feet. that is a linear measure. you could say it is 10 square feet

OpenStudy (anonymous):

the ^2 pertains the the letters, not the numbers

OpenStudy (anonymous):

yeah what ecin said

OpenStudy (anonymous):

follow this, Let x be the original side A=169 A=s^2 169=(x+8)^2, we can extract the square root of both sides 13=x+8 x=13-8 x=5 cm , this is the original side

OpenStudy (anonymous):

so u would say for 169cm^2 , 169 square feet. so u just squared the 169? then - from 8

OpenStudy (anonymous):

satellite wait till she got to volume

OpenStudy (anonymous):

cubic units

OpenStudy (anonymous):

if you had something that's 16 sq ft and it is square in shape

OpenStudy (anonymous):

no, square root of 169 is 13.

OpenStudy (anonymous):

then each side would be 4, .. and the area is 4^2

OpenStudy (anonymous):

actually i did not square the 169. i thought of a number whose square is 169

OpenStudy (anonymous):

sry pat that is what I meant not square feet

OpenStudy (anonymous):

in other words \[\sqrt{169}=13\]

OpenStudy (anonymous):

yeah, but that is having the same concept

OpenStudy (anonymous):

but that's only the case of a truly square shape, for now you should just ignore the sq units and just put it on the answer as a definition

OpenStudy (anonymous):

it is okay...

OpenStudy (anonymous):

l*w= A * some^2 unit

OpenStudy (anonymous):

yeah well i hope to not take anymore math after this class, keeps me stresses grrrr..lol

OpenStudy (anonymous):

quit while you are ahead!

OpenStudy (anonymous):

a cm * b cm = a* b cm^2, not cm^2 will always be there, so take it as a definition

OpenStudy (anonymous):

in my case, i do pay attention because its' not always the same. I'm in engineering

OpenStudy (anonymous):

just follow what i did awarr28

OpenStudy (anonymous):

oh ok ecin, at least I'm taking notes, but half the time I have to keep coming back to ask because i still cant figure it out...lol

OpenStudy (anonymous):

start with basic things, and things you can conceptualize

OpenStudy (anonymous):

yes I wrote that down pat, thank you all for all your help

OpenStudy (anonymous):

you are welcome^_^

OpenStudy (anonymous):

say you have tiled floors and they are all 1 feet

OpenStudy (anonymous):

or mess around with the little squares in graph papers think about them as square feets

OpenStudy (anonymous):

ok, I thank you all for your help.

OpenStudy (anonymous):

1 feet by 2 feet = 2 squares

OpenStudy (anonymous):

its like you are shading every little box or tile you count

OpenStudy (anonymous):

oh ok I see what your saying

OpenStudy (anonymous):

2 feet by 3 feet are six boxes

OpenStudy (anonymous):

and that's the area

OpenStudy (anonymous):

dont' worry about why, but just put it at the end ft^2

OpenStudy (anonymous):

so 2 feet by 6 feet would be 12 boxes right?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

;)

OpenStudy (anonymous):

you can also think about it like this "clear the area"

OpenStudy (anonymous):

you don't want nothing or nobody in that space

OpenStudy (anonymous):

a perimiter on the other hand is just like a gate.. is a measurement of the outside

OpenStudy (anonymous):

gotcha